1. Draw the quark flow diagrams for the following decays: (a) B_{s}^{0} → J/ψϕ,

(b) D^{−} → K^{0}π^{−}.

{4}

2. Draw one Feynmann diagram for each of these: (a) the decay of a top quark,

(b) the decay of a Higgs to four muons.

{4}